A rope is used to pull a 2.89-kg bucket of water out of a deep well. What is the acceleration of the bucket when the tension in the rope is 30.2 N?

Grade:11

2 Answers

Vikas TU
7 years ago
T – mg = ma
a = (T-mg)/m
  = (30.2 – 28.9)/2.89 
  = 0.5 m/s^2 approx. would be the accln. of the rope required to pull the 2.89 kg bucket of water.
Mars
6 years ago
\sum F= T - W ma = T - mg 2.89(a) = 30.2 - 2.89(9.81) a = 0.64 m/s